5x=3(1-y)
3x+3y+1=0

Answer by stanbon(75887)

5x=3(1-y)
3x+3y+1=0
-----
Rearrange::
5x + 3y = 3
3x + 3y = -1
--------
Subtract and solve for "x":
2x = 4
x = 2
-------------
Solve for "y"::
3x + 3y = -1
6 + 3y = -1
3y = -7
y = -7/3
-------------
Checking the answer::
5x=3(1-y)
3x+3y+1=0
Checking (2,-7/3)
-------
5(2) = 3(1+(7/3)
10 = 3+7
10 = 10
----
3(2) + 3(-7/3) + 1 = 0
6 - 7 + 1 = 0
0 = 0
---------
Cheers,
Stan H.
----------